OpenStudy (anonymous):

400 cm3 of mercury at 0 degrees Celcius will expand to what volume at 50 degrees Celcius?

OpenStudy (anonymous):

403.6 cm^3

OpenStudy (anonymous):

How did you get that? Please explain

OpenStudy (anonymous):

If the volume V0 of mercury is heated from t1 degree C to t2 degree C, then the volume v1 of mercury after expansion is given by: v1 = v0(1+gamma(t2-t1)), where gamma is the coefficient of volume expansion of mercury and is given to be 180*10^-6 /C. Substituting the given values, v1= 400(1+180*10^-6*(50-0)) =403.6 cm^3.

OpenStudy (anonymous):

So your equation is \[V _{1} = V _{0} ( 1 + 180 \times 10^{-6}) (50 - 0)\]

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

So, if you go through and solve you should get \[403.6 cm^{2}\]

OpenStudy (anonymous):

I got it. Thank you very much! Mind helping me with a few more questions?

OpenStudy (anonymous):

403.6cm^3*

OpenStudy (anonymous):

I'll answer if I can help. Don't forget to award me best response and close this question. ;P